OpenStudy (anonymous):

2y = x + 3 5y = x - 7 What is the solution set of the given system? {(-1/3, -4/3)} {(1/3, 4/3)} {(-29/3, -10/3)} {(29/3, 10/3)}

OpenStudy (pradipgr817):

multiply second equation with -1 and add to 1st -3y=10 y=-10/3 x=2y-3=(-20/3)-3=-29/3